# What Makes a Good Doctor Response? An Analysis on a Romanian Telemedicine Platform

### v1

@@ 5.  Conclusions
  • A. Cosma, O. Szehr, D. Kletz, A. Antonucci, and O. Pelletier (2026) Automatic prompt optimization for dataset-level feature discovery. External Links: 2601.13922, Link Cited by: §5.
@@
  • A. Niculae, A. Cosma, C. Dumitrache, and E. Radoi (2025a) Dr. copilot: a multi-agent prompt optimized assistant for improving patient-doctor communication in Romanian. In Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing: Industry Track, S. Potdar, L. Rojas-Barahona, and S. Montella (Eds.), Suzhou (China), pp. 1780–1792. External Links: Link, Document, ISBN 979-8-89176-333-3 Cited by: §2.
  • A. Niculae, A. Cosma, and E. Radoi (2025b) A retrieval-based approach to medical procedure matching in romanian. arXiv preprint arXiv:2503.20556. Cited by: §1.
@@
  • A. Rogoz, R. T. Ionescu, A. Anghel, I. Antone-Iordache, S. Coniac, and A. I. Ionescu (2025) MedQARo: a large-scale benchmark for evaluating large language models on medical question answering in romanian. arXiv preprint arXiv:2508.16390. Cited by: §1.
@@
  • D. Tufiș and V. Barbu Mititelu (2014) The lexical ontology for romanian. In Language Production, Cognition, and the Lexicon, N. B. Nuria Gala (Ed.), Text, Speech and Language Technology, Vol. 48, pp. 491–504. Cited by: §1.

### v3

@@ 6.  Bibliographical References
  • J. Bekker and J. Davis (2020) Learning from positive and unlabeled data: a survey. Machine learning 109 (4), pp. 719–760. Cited by: §3.
  • A. Cosma, O. Szehr, D. Kletz, A. Antonucci, and O. Pelletier (2026) Automatic Prompt Optimization for Dataset-Level Feature Discovery. External Links: Link, 2601.13922 Cited by: §5, §5.
@@
  • A. Niculae, A. Cosma, C. Dumitrache, and E. Radoi (2025a) Dr. Copilot: A Multi-Agent Prompt Optimized Assistant for Improving Patient-Doctor Communication in Romanian. In Proceedings of the 2025 Conference on Empirical Methods in Natural Language Processing: Industry Track, S. Potdar, L. Rojas-Barahona, and S. Montella (Eds.), Suzhou (China), pp. 1780–1792. External Links: Document, ISBN 979-8-89176-333-3, Link Cited by: §2.
  • A. Niculae, A. Cosma, and E. Radoi (2025b) A Retrieval-Based Approach to Medical Procedure Matching in Romanian. In Proceedings of the 24th Workshop on Biomedical Language Processing, D. Demner-Fushman, S. Ananiadou, M. Miwa, and J. Tsujii (Eds.), Viena, Austria, pp. 167–175. External Links: Document, ISBN 979-8-89176-275-6, Link Cited by: §1.
@@
  • J. W. Pennebaker, M. E. Francis, R. J. Booth, et al. (2001) Linguistic inquiry and word count: liwc 2001. Mahway: Lawrence Erlbaum Associates 71 (2001), pp. 2001. Cited by: §1.
@@
  • A. Rogoz, R. T. Ionescu, A. Anghel, I. Antone-Iordache, S. Coniac, and A. I. Ionescu (2026) A large-scale benchmark for evaluating large language models on medical question answering in Romanian. npj Digital Medicine. External Links: Document, ISBN 2398-6352, Link Cited by: §1.
@@
  • D. Tufiș and V. Barbu Mititelu (2014) The Lexical Ontology for Romanian. In Language Production, Cognition, and the Lexicon, Text, Speech and Language Technology, Vol. 48, pp. 491–504. Cited by: §1.